Adjustments Kretz Corporation prepares monthly financial statements and therefore adjusts its accounts at the end of every moHow does this entry affect the accounting equation? If a financial statement item is not affected, select No Entry and leavHow does this entry affect the accounting equation? If a financial statement item is not affected, select No Entry and leavHow does this entry affect the accounting equation? If a financial statement item is not affected, select No Entry and leav
Answer #1

Section 1

Part A

Activity

Operating

Accounts

Interest expense, interest payable

Statement (s)

Income statement and balance sheet

Balance Sheet

Income statement

Assets

=

Liabilities

+

Stockholders’ equity

Revenue

-

Expenses

=

Net income

interest payable

100

-100

interest expense

100

-100

(15000*8%*30/360 = 100)

Part B

Activity

Operating

Accounts

supplies expense, supplies

Statement (s)

Income statement and balance sheet

Balance Sheet

Income statement

Assets

=

Liabilities

+

Stockholders’ equity

Revenue

-

Expenses

=

Net income

supplies

-660

-660

Supplies expense

660

-660

1280+750-1370 = 660

Part C

Activity

Operating

Accounts

Depreciation expense, accumulated depreciation –office equipment

Statement (s)

Income statement and balance sheet

Balance Sheet

Income statement

Assets

=

Liabilities

+

Stockholders’ equity

Revenue

-

Expenses

=

Net income

Accumulate depreciation – office equipment

-800

-800

Depreciation expense

800

-800

($62,600 – $5,000) × 1/72 = 800

Part D

Activity

Operating

Accounts

Wages expense, Wages payable

Statement (s)

Income statement and balance sheet

Balance Sheet

Income statement

Assets

=

Liabilities

+

Stockholders’ equity

Revenue

-

Expenses

=

Net income

Wages payable

5700

-5700

Wages expense

-5700

-5700

950*6 = 5700

Part E

Activity

Operating

Accounts

Rent revenue, rent revenue collected in advance

Statement (s)

Income statement and balance sheet

Balance Sheet

Income statement

Assets

=

Liabilities

+

Stockholders’ equity

Revenue

-

Expenses

=

Net income

rent revenue collected in advance

-2500

2500

Rent revenue

2500

2500

Part F

Activity

Operating

Accounts

Customer Deposits, service revenue

Statement (s)

Income statement and balance sheet

Balance Sheet

Income statement

Assets

=

Liabilities

+

Stockholders’ equity

Revenue

-

Expenses

=

Net income

Customer Deposits

-1200

1200

Service revenue

1200

1200

Part G

Activity

Operating

Accounts

Income tax expense, income tax payable   

Statement (s)

Income statement and balance sheet

Balance Sheet

Income statement

Assets

=

Liabilities

+

Stockholders’ equity

Revenue

-

Expenses

=

Net income

Income tax payable   

3900

-3900

Income tax expense

3900

-3900

Section 2

Adjusted net income = reported net income – interest expense – supplies expense –depreciation expense - Wages expense + rent revenue + service revenue –income tax expense = 23000-100-660-800-5700+2500+1200-3900 = $15540